# The gravitational-wave emission from the explosion of a 15 solar mass star with rotation and magnetic fields

### v1

@@ 8 Conclusions
In recent years, there has been a significant increase in the number of long-duration 3D simulations of gravitational waves from typical CCSNe with no rotation and no magnetic fields. The gravitational-wave emission for neutrino-driven explosions still has some uncertainties, e.g., regarding signal amplitude, but the typical overall shape of the signal in the time-frequency space is well understood. However, there has only been a small number of gravitational-wave signals from 3D simulations that include magnetic fields and rotation that are continued well beyond the core-bounce phase. Therefore, in this paper, we performed four simulations of the explosions of a $15\,M_{\odot}$ star with weak pre-collapse magnetic fields to determine the impact of the magnetic fields on the gravitational-wave signal. We included both rotating and non-rotating models and used two different EoS.
@@
The end times of our simulations are between 0.42 s and 0.48 s after core bounce, which is long enough to capture the main f/g-mode and any gravitational-wave emission due to the SASI. We gave a brief discussion of the hydrodynamics that are relevant for understanding the gravitational-wave emission. The rotating models undergo shock revival at $\sim 100$ ms post bounce, and the non-rotating models undergo shock revival at $\sim 200$ ms post bounce. The observed explosion energies are similar to those observed in models with no magnetic fields, showing that our weak magnetic fields did not have a big impact on the explosion dynamics. The rotating models quickly form unstable jets.
@@
We also analysed the properties of the PNS, as the gravitational-wave frequency of the main f/g-mode is closely correlated with the PNS properties. We observe smaller PNS masses in rotating models, which is consistent with previous work. We have one model SFHx_rr that has a smaller PNS radius than all the other models. This results in the gravitational-wave emission for that model reaching higher frequencies than the other three models.
@@
The gravitational-wave amplitudes, and maximum detection distances are consistent with our previous models that do not include magnetic fields. Stronger magnetic fields would be needed for a CCSN to produce powerful gravitational-wave amplitudes from a magnetorotational explosion. The gravitational-wave frequencies are higher than in our previous neutrino-driven explosions, which is likely due to the simulation set up, but may also be partly due to the presence of the magnetic fields.
@@
We also investigate the low-frequency gravitational-wave emission due to both the asymmetric matter and the asymmetric emission of neutrinos. The low-frequency gravitational-wave emission from matter starts to grow to tens of cm in amplitude before the end of the simulation time, and would likely have reached a larger amplitude if we were able to continue the simulations for a longer duration. However, the gravitational waves due to the asymmetric emission of neutrinos have very high amplitudes, a few hundred cm at low-frequencies, by the end of the simulation time.
@@
We investigated the impact of the low-frequency gravitational-wave emission on the detectability of CCSNe, as this aspect of the gravitational-wave signal is neglected in most detectability studies. We show that the gravitational-wave amplitude at low-frequencies can vary significantly at different source angles. The gravitational waves from the asymmetric emission of neutrinos has no impact on the detectability of CCSNe in Advanced LIGO, as it occurs below the frequency band of current gravitational-wave detectors. However, we find that including this aspect of the gravitational-wave emission can have a significant impact on the detectability of CCSNe in the next generation detector Einstein Telescope, which shows that it is important that future studies include this feature of the gravitational-wave emission when developing the science case for detecting CCSNe in next generation detectors. We also show that CCSNe are a promising source for a moon-based gravitational-wave detector, even if we cannot predict our gravitational-wave frequencies low enough to cover the entire moon detection band, due to the short duration of our simulations.
@@ Acknowledgements
Authors JP and BM are supported by the Australian Research Council (ARC) Centre of Excellence (CoE) for Gravitational Wave Discovery (OzGrav) project numbers CE170100004 and CE230100016. JP is supported by the ARC Discovery Early Career Researcher Award (DECRA) project number DE210101050. BM acknowledges supported by the ARC through Future Fellowship FT160100035 and Discovery Project DP240101786. We acknowledge computer time allocations from Astronomy Australia Limited’s ASTAC scheme, the National Computational Merit Allocation Scheme (NCMAS), and from an Australasian Leadership Computing Grant. Some of this work was performed on the Gadi supercomputer with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government, and through support by an Australasian Leadership Computing Grant. Some of this work was performed on the OzSTAR national facility at Swinburne University of Technology. OzSTAR is funded by Swinburne University of Technology and the National Collaborative Research Infrastructure Strategy (NCRIS).

### v2

@@ 8 Conclusions
In recent years, there has been a significant increase in the number of long-duration 3D simulations of GWs from typical CCSNe with no rotation and no magnetic fields. The GW emission for neutrino-driven explosions still has some uncertainties, e.g., regarding signal amplitude, but the typical overall shape of the signal in the time-frequency space is well understood. However, there has only been a small number of GW signals from 3D simulations that include magnetic fields and rotation that are continued well beyond the core-bounce phase. Therefore, in this paper, we performed four simulations of the explosions of a $15\,M_{\odot}$ star with weak pre-collapse magnetic fields to determine the impact of the magnetic fields on the GW signal. We included both rotating and non-rotating models and used two different EoS.
@@
The end times of our simulations are between 0.42 s and 0.48 s after core bounce, which is long enough to capture the main f/g-mode and any GW emission due to the SASI. We gave a brief discussion of the hydrodynamics that are relevant for understanding the GW emission. The rotating models undergo shock revival at $\sim 100$ ms post bounce, and the non-rotating models undergo shock revival at $\sim 200$ ms post bounce. The observed explosion energies are similar to those observed in models with no magnetic fields, showing that our weak magnetic fields did not have a big impact on the explosion dynamics. The rotating models quickly form unstable jets.
@@
We also analysed the properties of the PNS, as the GW frequency of the main f/g-mode is closely correlated with the PNS properties. We observe smaller PNS masses in rotating models, which is consistent with previous work. We have one model SFHx_rr that has a smaller PNS radius than all the other models. This results in the GW emission for that model reaching higher frequencies than the other three models.
@@
The GW amplitudes, and maximum detection distances are consistent with our previous models that do not include magnetic fields. Stronger magnetic fields would be needed for a CCSN to produce powerful GW amplitudes from a magnetorotational explosion. The GW frequencies are higher than in our previous neutrino-driven explosions, which is likely due to the simulation set up, but may also be partly due to the presence of the magnetic fields.
@@
We also investigate the low-frequency GW emission due to both the asymmetric matter and the asymmetric emission of neutrinos. The low-frequency GW emission from matter starts to grow to tens of cm in amplitude before the end of the simulation time, and would likely have reached a larger amplitude if we were able to continue the simulations for a longer duration. However, the GWs due to the asymmetric emission of neutrinos have very high amplitudes, a few hundred cm at low-frequencies, by the end of the simulation time.
@@
We investigated the impact of the low-frequency GW emission on the detectability of CCSNe, as this aspect of the GW signal is neglected in most detectability studies. We show that the GW amplitude at low-frequencies can vary significantly at different source angles. The GWs from the asymmetric emission of neutrinos has no impact on the detectability of CCSNe in Advanced LIGO, as it occurs below the frequency band of current GW detectors. However, we find that including this aspect of the GW emission can have a significant impact on the detectability of CCSNe in the next generation detector Einstein Telescope, which shows that it is important that future studies include this feature of the GW emission when developing the science case for detecting CCSNe in next generation detectors. We also show that CCSNe are a promising source for a moon-based GW detector, even if we cannot predict our GW frequencies low enough to cover the entire moon detection band, due to the short duration of our simulations.
@@ Acknowledgements
Authors JP and BM are supported by the Australian Research Council (ARC) Centre of Excellence (CoE) for Gravitational Wave Discovery (OzGrav) project numbers CE170100004 (JP, BM) and CE230100016 (JP). JP is supported by the ARC Discovery Early Career Researcher Award (DECRA) project number DE210101050. BM acknowledges supported by the ARC through Future Fellowship FT160100035 and Discovery Project DP240101786. We acknowledge computer time allocations from Astronomy Australia Limited’s ASTAC scheme, the National Computational Merit Allocation Scheme (NCMAS), and from an Australasian Leadership Computing Grant. Some of this work was performed on the Gadi supercomputer with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government, and through support by an Australasian Leadership Computing Grant. Some of this work was performed on the OzSTAR national facility at Swinburne University of Technology. OzSTAR is funded by Swinburne University of Technology and the National Collaborative Research Infrastructure Strategy (NCRIS).